# What Matters in Hierarchical Search for Combinatorial Reasoning Problems?

### v1

@@ 4.1.1 Verifying experimental outcomes
Low-level search algorithms are competitive with subgoal methods for solving the Rubik’s Cube
4.1.2 Learning from diverse experts
Motivation
Subgoal methods are considered more effective for long-horizon planning problems [7] and are believed to reduce noise in training data [14]. To verify these properties, we trained the algorithms to solve the Rubik’s Cube using a dataset composed of various sources: human-oriented Beginner and CFOP, computational Kociemba, and the random solutions used in Section 4.1.1. The first two solvers solve the cube layer by layer; hence, their trajectories are long and composed of high-level subroutines. The remaining two usually find much shorter solutions (see Appendix B.2 for examples).
Experimental Results
When training on data from diverse experts, the subgoal methods (particularly AdaSubS) outperform the low-level search by a wide margin (see Figure 2). Notably, despite using strongly suboptimal and diverse data, AdaSubS and kSubS can solve the most challenging instances within an even smaller budget than in the previous experiment. On the other hand, all low-level methods perform much worse than in the previous experiment. Additionally, we observe that given lengthy demonstrations, increasing the subgoal distance for kSubS improves performance, unlike in Section 4.1.1 with short trajectories.
We attribute the observed performance gap to the noise in the value function estimates. Specifically, it fails to consistently assign low values to fully scrambled cubes and nearby states. This issue arises because both Beginner and CFOP yield solutions that are even five times longer than those of the other solvers. Subgoal algorithms are, to some degree, robust to this noise, as claimed by [14]. By taking longer steps, they can leap over the faulty value function. A detailed analysis of this behavior is provided in Appendix B.2.
@@ 4.1.2 Learning from diverse experts
Subgoal methods successfully leverage diverse, long demonstrations, while low-level search performs better on short, homogeneous trajectories.
Figure 2: Solving the Rubik’s Cube. Components are trained on data from 4 different experts.
4.1.3 Structure in demonstrations
Motivation
While Section 4.1.2 showcases an advantage of subgoal methods, it is still unclear which property of the environment contributes the most to it. We conjecture that the key factor is the diversity of training data, which makes the value estimation problem hard. To verify this, we use training data collected from a single algorithmic solver – Beginner. Unlike in the previous experiment, no other expert is used. Therefore, the training data, although not optimal, becomes unambiguous.
Figure 4: Success rate of solving the Rubik’s Cube within a budget of $500$ visited states, with artificial Gaussian noise added to each value estimate.
Experimental Results
After removing the variance from demonstrations by switching to Beginner, the performance of AdaSubS and BestFS becomes nearly identical, as illustrated in Figure 4. Similarly, the performance of A* closely matches that of kSubS. MCTS shows significant improvement over previous experiments, though it still lags behind other methods. Longer subgoals continue to outperform shorter ones, but extending them beyond a distance of four offers no additional benefits. Furthermore, a comparison with the results in Section 4.1.2 reveals that the performance of subgoal methods remains consistent whether using single or multiple sources of demonstrations.
Importantly, we found that in order to achieve good performance, BestFS requires careful tuning. It achieves its best results when, at each expansion, we select actions that collectively exceed $70\%$ confidence. Additionally, we found that its performance is highly sensitive to this hyperparameter (see Appendix F). Adopting the approach recommended by [14, 63] of selecting a fixed number of actions leads to highly suboptimal outcomes, as demonstrated by the poorly performing agent in Figure 4, BestFS-top-2.
@@ 4.1.3 Structure in demonstrations
Given homogeneous data, low-level search algorithms can leverage the long, structured demonstrations equally well as subgoal methods.
4.1.4 Robustness to noisy value function
Motivation
The advantage of subgoal methods observed in Section 4.1.2 does not primarily come from long or structured demonstrations. However, high-level algorithms show surprisingly consistent performance regardless of the training data, as demonstrated in Sections 4.1.1-4.1.3. That observation hints that the subgoal methods are resilient to noise in the training data. To verify that proposition, we take the components trained as in Section 4.1.1 and add artificial Gaussian noise to each output of the value function.
Experimental Results
As shown in Figure 4, while the standard BestFS can solve nearly all instances, its performance heavily depends on the quality of the learned value function. In a similar way, the noisy values strongly impact the performance of A* and MCTS. However, kSubS is much more resilient when using subgoals longer than a single step. Even with completely random values ($\sigma=100$), it successfully solves about $40\%$ instances. Moreover, AdaSubS is barely affected by the noise at all. That confirms the claim of [14] that subgoals successfully leap over value noise.
As noted in [63], the search performed by the subgoal methods is strongly guided by the subgoal generators, which considerably reduces the reliance on the value function. A more detailed analysis of this experiment can be found in Appendix B.3.
@@ 4.1.4 Robustness to noisy value function
Subgoal methods successfully leap over the noise. Thus, they should be used when estimating the value is hard, for instance, when learning from diverse and suboptimal experts.
4.2 INT
4.2.1 Scaling to complex mathematical reasoning
Motivation
We aim to explore whether subgoal methods can more effectively scale to complex real-world applications than low-level methods, as suggested in [14, 63]. To this end, we consider the INT environment [60], which involves solving inequalities. It is the most challenging among the environments studied in [14, 63, 35, 36], both in terms of observations and actions complexity.
Experimental Results
Our experiments show a clear advantage of subgoal methods compared to low-level search on INT, as shown in Figure 6. We note that we use the proofs length of $25$, which is a more challenging setup than typically used.
@@ 4.2.1 Scaling to complex mathematical reasoning
Subgoal methods significantly outperform low-level search strategies in the context of mathematical reasoning.
4.2.2 Complex action space
Motivation
We conjecture that the result of the previous experiment is due to the high complexity of the action space compared to the state space. This suggests that state-based subgoals may be more effective in such settings than action-based policy guidance. To further investigate this in a controlled setting, we inflate the action space of the Rubik’s Cube by a factor of $100$ by creating $100$ copies of each original action.
Experimental Results
As shown in Figure 6, the low-level search methods stuck below $30\%$ success rate, while the subgoal methods perform comparably to the standard cube case from Section 4.1.1. We could expect that since, given so many similar actions, the low-level search nearly trivializes. Although the subgoal methods also use a low-level policy for connecting the subgoals, it always follows a single path towards the subgoal, so it is barely affected. A more detailed analysis can be found in Appendix B.4.
Takeaway
When facing a problem with a complex action space, subgoal methods should outperform low-level search.
Figure 5: Solving INT. Components are trained on randomly generated proofs of length $25$.
Figure 6: Solving the Rubik’s cube with inflated action space.
4.3 Sokoban
4.3.1 Importance of budget measures
Motivation
Given that Sokoban is PSPACE-complete [13], it presents a challenging testbed for AI planning algorithms. [63] reports that AdaSubS outperforms BestFS in Sokoban. They also report that kSubS surpasses BestFS only, given a relatively high search budget. Some works use the number of high-level nodes visited during the search as the search budget [35]. This contrasts with the total number of visited nodes (the main metric in this paper). We aim to show how this difference affects the interpretation of the results and propose recommendations for reporting the search budget. For these experiments, each method was trained on a dataset containing trajectories collected by an external RL agent.
Experimental Results
Figure 8 shows that reporting only the high-level nodes excessively enhances subgoal methods’ scores. AdaSubS and kSubS seem absolutely superior to BestFS despite the last two algorithms visiting a similar number of states during the search. Hence, that way of reporting the results would lead to erroneous conclusions about the strong primacy of subgoal methods, and we recommend using the total number of visited states as the search budget.
Takeaway
When reporting the search budget, it is crucial to count all visited states, not only the high-level nodes.
Search Algorithm Dead-Ends
MCTS 22.0%
BestFS-$85\%$ 18.5%
A* 13.7%
kSubS-$4$ 12.7%
kSubS-$8$ 10.0%
AdaSubS-$8+4$ 8.86%
Figure 7: Comparison of the fraction of dead-ends encountered during the problem-solving process between subgoal-based and low-level searches.
Figure 8: Comparison of two meanings of search budget: total states visited versus only high-level nodes visited. For BestFS, both definitions are equivalent.
4.3.2 Dead-ends
Motivation
We call a state a dead-end if it is impossible to reach the goal from it. They pose an additional challenge as an algorithm must quickly recognize dead-ends to abandon the search in hopeless state-space regions. We investigate how the subgoal and low-level methods handle dead-ends.
Experimental Results
Table 8 indicates that subgoal methods, especially those involving longer subgoals, notably reduce the likelihood of stacking in dead-ends. We conjecture that this effectiveness stems mainly from relying less on value function guidance than low-level methods, as detailed in Section 4.1.4 and supported by [20]. For further analysis, refer to Appendix B.1.
Takeaway
Subgoal methods are better in avoiding dead-ends than low-level search.
4.4 N-Puzzle
Motivation
The N-puzzle, also known as the Sliding Tile Puzzle, is a classic benchmark in the search literature [41, 35, 34] and is known to be NP-hard [46]. It was used to evaluate subgoal methods in [35, 36]. Despite its large state space, the N-puzzle is highly structured, allowing for solving very simple suboptimal strategies. We aim to determine whether subgoal techniques offer an advantage over low-level search methods in this environment.
Experimental Results
To generate the training dataset for this experiment, we employed a sub-optimal algorithm that orders the tiles in ascending order, starting with "$1$" and continuing through "$2$", "$3$", and so on (details in Appendix A).
Figure 9: Solving the N-Puzzle. The success rates of BestFS, A*, and BC policy throughout all budgets match nearly perfectly.
A simple policy trained with behavioral cloning (BC) perfectly solves the environment, even without using search, see Figure 9. This indicates that learning from structured demonstrations is relatively straightforward. Notably, AdaSubS surpassed BC, indicating that this subgoal approach improved the cloned sub-optimal behavior. Furthermore, kSubS, BestFS, and A* perform similarly to the BC policy, showcasing no computational waste.
Takeaway
Due to its structure, the N-Puzzle environment is not a good benchmark for testing the performance of advanced search algorithms.
@@ 5 Open questions and future directions
While we identified several features that facilitate the performance of subgoal methods, that list is not exhaustive. Thus, it is essential to study this topic further, expand the analysis to more subgoal-based and low-level algorithms, and include even more types of environments. Additionally, our work provides mostly experimental validation of the claims. Finding theoretical foundations for the observed properties would also be a valuable direction.
@@ 6 Conclusions
We conducted a thorough comparison of subgoal-based and low-level search algorithms for combinatorial reasoning tasks. Our experiments indicate that hierarchical approaches excel in environments where value estimation is challenging and learned estimates face significant uncertainty, particularly when learning from various suboptimal experts. Furthermore, subgoal methods demonstrate better scalability in complex action spaces and are more effective at avoiding dead ends than low-level methods. Thus, in environments characterized by those properties, it is advisable to consider subgoal-based methods.
According to our experiments, the best-performing low-level search was usually BestFS with a confidence threshold. Although it is rather sensitive to the threshold value, which has to be optimized for each domain separately, we advocate using it as a standard baseline for further research in hierarchical search.
Finally, we identified a few easy-to-overlook mistakes in reporting the results that may lead to misleading conclusions. Most importantly, the reported search budget of the hierarchical methods must include all the visited states and not only the high-level nodes as used in some prior works.
@@ Appendix A Environments
INT INT (INequality Theorem proving) is an automated theorem-proving benchmark for high school algebraic inequality proofs. [60] provides a generator of mathematical inequalities and a proof verification tool. Each action in INT maps to a proof step, which specifies a chosen axiom and its input entities - which makes action space very high-dimensional, enabling up to a million valid actions at a step. This large action space makes INT a desirable but challenging environment for expanding HRL paradigms to vast action spaces.
We used 25-step proofs for this paper, representing an uplift from 15 considered in [14, 63] (the latter used longer proofs, but only for evaluating 15-trained models). Each step is an application of an axiom to an axiom-specific number of entities (entities are bracketed or bracketable parts of the theorem’s goal).

### v3

@@ 5.1 Subgoal Methods Benefit from Diverse Sources of Data
Subgoal methods successfully leverage diverse demonstrations (Q2), while low-level search performs better when trained on homogeneous trajectories (Q2).
5.2 Subgoal Methods are Value Noise Filters
We found that the classical search algorithms are highly sensitive to the quality of the value function. To show this in a controlled setting, we added Gaussian noise to the value estimates and observed how different noise levels impacted the success rate of solving tasks.
Figure 7: Success rate of low-level and subgoal methods as the approximation errors of the value function increase. $\sigma=100$ results in completely random value estimates.
While $\rho$-BestFS is able to solve nearly all instances under ideal conditions, its performance significantly declines as value function errors increase, even to $0\%$ (see Figure 7). $\rho$-A* and $\rho$-MCTS behave similarly. In contrast, the subgoal methods show remarkable resilience. Particularly AdaSubS, which maintains nearly unchanged success rate, despite high value errors (Q2).
These results align with our findings in Section 5.1, where using diverse training data naturally introduced value estimation errors. As observed by Zawalski et al. (2023), the search process of subgoal methods is guided by subgoal generators, which reduces reliance on the value function. Subgoal generators and the conditional policies connecting subgoals are not directly influenced by the value approximation errors. The value function is used only in high-level nodes, which represent only a fraction of the search tree.
In hierarchical methods, the distance between high-level nodes spans multiple steps, increasing the likelihood that value estimates for subsequent high-level nodes along the solution path will be monotonic (see Figure 9 for an illustrative example), which makes planning more efficient. This supports the claim by Czechowski et al. (2021) that subgoals effectively mitigate the impact of value noise. To further ground that result, we prove the following theorem:
Theorem 1 (Search advancement formula).
Let $g_{k}:S\to\mathcal{P}(S)$ be a stochastic $k$-subgoal generator that, given a state $s\in S$ samples a set of $b$ subgoals $\{s_{i}\}$ such that the distances $d(s_{i},s)$ are independent, uniformly distributed in the interval $[-k;k]$. Let $V:S\to\mathbb{R}$ be a value function with approximation error uniformly distributed in the interval $[-\sigma;\sigma]$.
Then, after $n$ iterations of search, the expected total progress toward the goal is:
$\mathbb{E}_{Adv}=\frac{nb}{4\sigma k}\int_{-k}^{k}x\left(\int_{-\sigma}^{\sigma}\tilde{u}(x+h)^{b-1}\mathrm{d}h\right)\mathrm{d}x,$ (1)
where $\tilde{u}(x)$ is CDF of the sum of two uniform variables $U(-k,k)+U(-\sigma,\sigma)$. Additionally, if we approximate that sum as $U(-k-\sigma,k+\sigma)$, we get
$\mathbb{E}_{Adv}\approx\frac{n\left((k+\sigma)^{b}(bk^{2}+bk\sigma-2k\sigma-2\sigma^{2})+\sigma^{b}(2k\sigma+bk\sigma+2\sigma^{2})-k^{b}(bk^{2})\right)}{(b+1)(b+2)k\sigma(k+\sigma)^{b-1}}$ (2)
Proof.
See Appendix K for the proof. ∎
Figure 8: Value estimates along a solving trajectory generated by $\rho$-BestFS. Even small approximation errors cause non-decreasing values, slowing down the search. In contrast, the subgoal path mitigates these errors, leading to mostly monotonic values along the trajectory.
Figure 9: Normalized advancement ${\mathbb{E}_{Adv}}/{k}$ for a single search iteration, according to Theorem 1. The value for each subgoal is divided by its length to represent the advancement per atomic action for easier comparison.
Theorem 1 quantifies the expected progress of the search at each step, with Equation 1 giving an exact formula and Equation 2 providing a useful approximation. To compare subgoal methods with low-level methods in theory, under different levels of value approximation error, we model low-level search by setting $k=1$, which represents a single action. Figure 9 shows the expected search progress with a branching factor of $b=3$, normalized by the number of actions leading to a subgoal.
When value estimates are perfect (i.e., $\sigma=0$), both subgoal and low-level searches perform similarly. However, as value approximation errors increase, subgoal methods become significantly more resilient. At high noise levels ($\sigma=20$), single-step searches make very little progress, advancing only $0.025$ per action. In contrast, subgoals of length 8 achieve much greater progress – $1.4$ for the entire subgoal, which is $0.175$ per action. This 7-fold increase in theoretical efficiency explains why subgoal methods outperform low-level methods in our experiments.
Further analysis of these experiments can be found in Appendix B.2.
@@ 5.2 Subgoal Methods are Value Noise Filters
Subgoal methods successfully handle value approximation errors. Thus, they should be used when estimating the value is hard, for instance, when learning from diverse and suboptimal demonstrations (Q2).
5.3 Subgoal Methods Handle Complex Action Spaces
Figure 10: Solving INT. Components are trained on randomly generated proofs.
Figure 11: Solving the Rubik’s cube with expanded action space, compared with the standard setup. Components are trained on reverse random shuffles.
In environments with large action spaces, search methods often struggle due to the exponential increase in the number of choices (Sutton & Barto, 1998). As shown in Figure 11, subgoal methods demonstrate a clear advantage over low-level search methods in the INT environment (Wu et al., 2021), a benchmark on proving mathematical inequalities (Q1). The INT environment is particularly challenging because of its highly complex observation and action spaces, making it the most difficult benchmark among those used in (Czechowski et al., 2021; Zawalski et al., 2023; Kujanpää et al., 2023a; b).
Given a complex action space, each node expansion in low-level methods involves executing many similar actions, limiting their ability to efficiently search through the space. In contrast, subgoal methods compute actions only to connect subgoals, which is a much simpler task. This targeted approach reduces the negative impact of a large action space, allowing subgoal methods to maintain their efficiency even as the action space grows (Q2).
To confirm this explanation, we conducted experiments on a modified version of the Rubik’s Cube, where the action space was artificially inflated by giving the agent access to 100 copies of each action. As shown in Figure 11, this simple modification drastically reduces the success rates of all low-level methods, even below 35%. In contrast, subgoal methods remain largely unaffected, performing similarly to the standard setup. We can explain that result with the following theorem:
Theorem 2 (Densification of the action space).
Fix any state $s$ from the state space $S$. Let $f:A\to[0,1]$ be the action distribution induced by the data-collecting policy for the state $s$. Assume that $f$ is continuous and has a unique maximum.
For clarity, assume $A=[0,1]$. Consider a sequence of increasingly dense discrete action spaces $A_{n}:=\{i/n\}_{i=0}^{n}\subset A$. Let $\rho_{n}:S\times A_{n}\to[0,1]$ be a family of policies that learn the distribution $f|_{A_{n}}$ over actions, with uniform approximation error $U(-E,E)$, where $E\in\mathbb{R}_{+}$. Let $r_{n}$ be the range of the top $K$ actions according to the probabilities estimated by $\rho_{n}$. Then
$\lim_{n\to\infty}\mathbb{E}[r_{n}]=0.$
Proof.
See Appendix L
Intuitively, this theorem states that as the action space become more dense and complex, the actions sampled for search become increasingly less diverse, which strongly impedes successful planning. Note that this analysis is strictly more general than the last experiment, where we simply copied the available actions. Further analysis of the experiments involving large action spaces is provided in Appendix B.3.
@@ 5.3 Subgoal Methods Handle Complex Action Spaces
When facing a problem with a complex action space, subgoal methods should outperform low-level search (Q2).
5.4 Subgoal Methods Avoid Dead Ends
Search algorithm Dead ends rate
$\rho$-MCTS 22.0%
$\rho$-BestFS 18.5%
$\rho$-A* 13.7%
kSubS ($4$ steps) 12.7%
kSubS ($8$ steps) 10.0%
AdaSubS 8.86%
Figure 12: Fraction of dead ends encountered during search between hierarchical and low-level methods in Sokoban.
Once an agent encounters a dead end, reaching the goal becomes impossible, leading to wasted computational effort. Our results, presented in Figure 12, show that subgoal methods tend to enter dead ends less often than low-level methods. Using longer subgoals improves the ability to bypass those areas.
Among low-level methods, $\rho$-A* performs the best at minimizing dead ends rate, as its node selection regularizes values by depth in the search tree, preventing it from over-committing to dead ends. However, even $\rho$-A* is outperformed by subgoal methods, which rely on greedy value estimates and subgoals.
Deciding whether a state is a dead end can be NP-hard. Hence, it is much harder for the value function to penalize dead ends compared to the policy, which only ranks the available actions and does not have to identify dead ends (Feng et al., 2022). Furthermore, demonstrations used for imitation learning lead to the goal state, hence they contain no dead ends. Therefore the value function trained this way is never directly instructed to penalize dead ends. At the same time, during training of the policy the actions leading to dead ends are never reinforced. Our experiments show that hierarchical search relies much less on the value guidance compared to low-level search (Section 5.2), which further supports our conclusions. For a more detailed analysis, see Appendix B.4.
@@ 5.4 Subgoal Methods Avoid Dead Ends
Subgoal Methods Are More Effective at Avoiding Dead Ends Compared to Low-Level Search (Q2).
5.5 Subgoal Methods Generalize Out-Of-Distribution
Planners that can generalize to out-of-distribution (OOD) instances are essential for robust decision-making (Kirk et al., 2023; Shen et al., 2021). We tested two types of generalization in the Sokoban environment: by significantly changing the layout of the board and by using extremely difficult boards from the DeepMind dataset (Guez et al., 2018) (see Figure 13 for examples).
Figure 13: Examples of Sokoban boards used in OOD experiments
Figure 14: Averaged OOD results on Sokoban boards with OOD layouts. These instances were generated by systematically varying all parameters of the instance generator.
Figure 15: Performance on DeepMind extra hard boards.
In both cases, subgoal methods show better performance than low-level methods, with the gap increasing as the distribution shift become more visible (see Figures 15-15). However, we found that kSubS, when using twice longer subgoals, collapses in OOD evaluations, despite outperforming $\rho$-BestFS and other low-level methods on in-distribution tasks. As the subgoal distance increases, predicting the distant future becomes more challenging, making it less likely for the generated subgoals to be valid and reachable, especially in OOD tasks. In contrast, low-level methods avoid this issue, as selecting an action from a limited set always results in a valid move. Thus, while subgoal methods can be effective in OOD scenarios, excessively long subgoals can degrade performance (Q2).
When evaluated on extremely challenging instances (see Figure 1) introduced by (Guez et al., 2018), all methods required a significantly higher search budget but maintained the same performance order as in the previous experiment (Q1). Solving these instances requires more advanced strategies than those learned during training. Subgoal methods are better equipped to handle this increased complexity because selecting subgoals is closely related to choosing a broader strategy because of their longer horizon. In contrast, low-level methods must assess each individual action, which limits their ability to foresee the long-term consequences of their choices.
@@ 5.5 Subgoal Methods Generalize Out-Of-Distribution
Subgoal methods can scale better than low-level methods on OOD instances, provided the subgoals are not too long (Q2).
@@ 6 Open Questions and Future Directions
While we identified several features that facilitate the performance of subgoal methods, that list is not exhaustive. Thus, it is essential to study this topic further, expand the analysis to more subgoal-based and low-level algorithms, and include even more types of environments. While most of our takeaways were confirmed in multiple environments, extending the evaluation to more domains would strengthen our conclusions. Additionally, our work provides mostly experimental validation of the claims. Finding theoretical foundations for the observed properties, such as Theorem 1, would also be a valuable direction.
@@ 7 Conclusions
We conducted a thorough comparison of hierarchical and low-level search methods for combinatorial reasoning tasks. Our experiments provides empirical and some theoretical evidence that hierarchical approaches should be preferred in environments where value estimation is challenging and learned estimates face significant uncertainty, particularly when learning from diverse suboptimal data. Furthermore, subgoal methods demonstrate better scalability in complex action spaces and are more effective at avoiding dead ends than low-level methods. Thus, in environments characterized by those properties, it is advisable to consider subgoal methods as an alternative to low-level search.
Based on our results, we propose guidelines for future research in this area. According to our experiments, the best-performing low-level search was usually $\rho$-BestFS with a confidence threshold (see Appendix F). Although it is rather sensitive to the threshold value, which has to be optimized for each domain separately, we advocate using this simple method as a standard baseline for further research in hierarchical search. Our guidelines are comprehensively discussed in Appendix J.
Additionally, we identified easy-to-overlook mistakes in reporting the results that may lead to misleading conclusions. Most importantly, the reported complete search budget of hierarchical methods must include all the visited states and not only the high-level nodes as used in some prior works.
8 Broader Impact
Our study has broader implications for other complex domains. For example, advancements in robotics often face significant challenges due to limited data, leading many methods to rely on collective datasets like Open X-Embodiment (Collaboration et al., 2023). As shown in our experiments, hierarchical search methods benefit substantially from training on diverse expert data (Section 5.1). Furthermore, the data bottleneck increases the need for the models to generalize to out-of-distribution scenes and tasks, which is also an advantage of hierarchical methods (Section 5.5). Finally, an essential aspect of robotics involves preventing the robot from becoming stuck or losing the manipulated object, events that can be seen as dead-end scenarios (Section 5.4). Successful applications of hierarchical methods in robotics include models such as SuSIE (Black et al., 2024) and HIQL (Park et al., 2023).
Additionally, our experiments indicate that hierarchical methods scale well in long-horizon tasks, as evidenced by their performance in the N-Puzzle and the Rubik’s Cube (using Beginner demonstrations), where the average sequence of steps often exceeds 200. Interestingly, while low-level methods can still perform well in these scenarios, we observed that they tend to be much more sensitive to hyperparameter tuning.
It is important to note that we do not claim hierarchical methods are universally superior to low-level approaches in all complex domains. Instead, the properties highlighted in our analysis suggest cases where they should be considered.
9 Reproducibility Statement
The code used to run all our experiments is available at https://github.com/subgoalsearchmatters/what-matters-in-hierarchical-search. We also link there datasets used for training our models. Hence, all our results are fully reproducible.
Appendix
[…]
@@ Appendix A Environments
INT INT (INequality Theorem proving) is an automated theorem-proving benchmark for high school algebraic inequality proofs. (Wu et al., 2021) provides a generator of mathematical inequalities and a proof verification tool. Each action in INT maps to a proof step, which specifies a chosen axiom and its input entities - which makes action space very high-dimensional, enabling up to a million valid actions at a step. This large action space makes INT a desirable but challenging environment for expanding HRL paradigms to vast action spaces.
We used 25-step proofs for this paper, representing an uplift from 15 considered in (Czechowski et al., 2021; Zawalski et al., 2023) (the latter used longer proofs, but only for evaluating 15-trained models). Each step is an application of an axiom to an axiom-specific number of entities (entities are bracketed or bracketable parts of the theorem’s goal).